ConfigTHOR

Configuration program for THOR

Release for THOR 2.6

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994-1998 Ultima Thule Software Edition 1 - 8 January 1999

Petter Nilsen Eivind Nordseth Kjell Irgens Troels Walsted-Hansen Borge Nost Eirik Nicolai Synnes Marius Mortensen

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1993-1998 Ultima Thule Software, All Rights Reserved. Published by Ultima Thule Software. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies and that no information is removed or added without Ultima Thule Software’s permission. Typeset in TEX.

1 About

ConfigTHOR is the configuration program for THOR. It offers a wide range of options to customize THOR for your needs.

2 Systems Configuration

This is the window used to configure new systems. You will have to enter some information:

The name of the system (BBS or Internet provider) you want to add.

The packet name the system uses. Remember to omit extensions. Note that a packet name is not mandatory and is only required on system types where you actually receive you mail in packet/file. TCP systems for Internet usage do not required a packet name, but if you choose to provide on, Thor will assume this is a mailbox and parse it as one. This can be handy for SMTPd where you receive your mail via SMTP instead of POP3. Please refer to the THOR.faq for how to set this up.

Some examples:
System Type:   Example grab:    Packet name:
------------   -------------    ------------
ABBS/MBBS      thule.lzx        thule
QWK            thule1256.qwk    thule
SOUP           soup3683.soup    soup
Omen           omenb1.zip       omenb1
BlueWave       romeo.zip        romeo
Hippo          girls.lzh        girls
Fido           1CA9B792.SU0     #?.((MO|TU|WE|TH|FR|SA|SU)[0-9]|PKT)
UUCP           john             john
Use the maximum number of letters possible to identify  the
grabs,  e.g.,  "WHT" if grabs are named "WHT010993.QWK" and the
numbers change from session to session.
Do not use the same packet name on two systems!

The type of system or message format on the system you are calling (choose from the list).

A few other configuration windows are used for other parameters on a system wide level:

User information

In this window, information about yourself can be entered. Not all of these options are used by all system types, and some of this information is collected from your keyfile (if you have registered THOR, which we hope you have).

Systems and Conference Configuration:
EMail Address

This is the electronic address you have on this system/conference. The EMail address is mostly used on Fido-systems and Internet-systems (SOUP and UUCP). So, if you use an Internet system type (TCP, TCP ONLINE, SOUP or UUCP), enter you email address here, if you use a Fido system type, enter you netmail address here.

Only Systems Configuration:
User Name

This is your real name. This name will be used as default on systems supported with a NComm script, and as your real name on Internet type of systems (like TCP, SOUP and UUCP).

Street Address

This is your street address.

Address

This is your city and zip code.

Country

This is the country you live in.

Phone

This is your phone number.

Files and paths

Tag file

This is the tag file. The file must have one tag on each line. A random tag from this file will be a added under the signature in messages you write. If you want to use multiline tags, this can be done by using backslash n (\n) to separate lines in the tag. The field can be left empty, in which case no tag will be used. A cycle gadget will let you choose to use the global defined tag file, don’t use a tag file on this system/conference or use the tag file you define here.

Systems and Conference Configuration:
Signature

This is the signature that is appended at the end of the message file before it is imported into the editor (and in front of any taglines). There are 3 ways of adding signatures:

String

The contents of the signature string gadget will contain the string to be imported into the message.

File

The contents of the signature string gadget is a file name of a text file that will be imported into the message.

Command

The contents of the signature string gadgets is a shell command that will receive arguments with the template "BB-SNAME/K,CONFNAME/K" when run and must write the output to stdout.

A cycle gadget lets you choose to use the global defined signature, don’t use a signature on this system/conference or use the signature you define here.

Entering Confs

If this window is opened from the Systems Configuration, this is the Arexx script that will be run whenever you enter any conference on this system. When in Conference Configuration, this is the Arexx script that will be run whenever you enter this conference. (See [ARexx scripts], page 11 for further information).

Leaving Confs

If this window is opened from the Systems Configuration, this is the Arexx script that will be run whenever you leave a conference on this system. When in Conference Configuration, this is the Arexx script that will be run whenever you leave this conference. (See [ARexx scripts], page 11 for further information).

Only Systems Configuration:
Download Dir

This is the download directory where grabs and packets that you download from systems will be searched for. This setting will override the globally defined download directory. This is not required for TCP systems, but if you provide one, it should usually be UUMAIL: provided you get you mail by some other means than POP3.

Upload Dir

This is the upload directory where reply packets will be placed. This setting will override the globally defined upload directory.

Entering System

This Arexx script will be run whenever you enter this system in THOR. (See [ARexx scripts], page 11 for further information).

Leaving System

This Arexx script will be run whenever you leave this system in THOR. (See [ARexx scripts], page 11 for further information).

Import and Export options

Only Systems Configuration:

Event Arc

This defines the archiver type to use when archiving the reply package. The default is the archiver defined in the system type, and you normally don’t have to change it.

Script Flags

If a system type is supported with a script, all or some of these flags will be available.

Get Newfiles

Request a list of new files from the remote system.

Auto DL Private

Automatically download files uploaded private to you.

Auto Log Off

Automatically log off after the script session has ended.

Get Bulletins

Automatically collect updated bulletins since you last logon.

Ansi Menus

Configure the remote system to use ANSI menus.

Use Colors

Configure the remote system to display colors while you are online.

Add users

If this gadget is checked, all users that receive or send messages will be added to the user database. If this gadget is NOT checked, only users in mail conferences will be added. On Internet type of systems, this option will only add users in email when it is selected. You should not use this option on systems with a lot of users, e.g. Internet news and Fidonet. Parsing will be slower if the user database is large.

7-bit Mail

This option applies on system types where email is sent using the Internet. Selecting this option will make sure all email is sent as 7-bit, in compliance with RFC 822. If the mail contains 8-bit characters, the mail will be encoded using MIME quoted-printable encoding.

Systems and Conference Configuration:
Charset

All messages in the message database is kept in ISO character set. The messages in converted from the default charset to ISO when imported. By setting this option to something else than "Default", you can select what charset the messages is to be converted from. Usually, the default setting will be correct.

Xpk Method

This is the Xpk method that will be used when packing messages, e.g. "NUKE" or "RAKE". The cycle gadget to the right can be set to use this setting, don’t use Xpk on this system/conference or to use the global settings (default).

Show Own Msgs

If this option is checked, messages written by you are marked as unread when imported into the database.

Line Length

By setting the contents of this integer gadget to something else than 0, you can define a custom line length to use when replying to messages. The default setting is 0, in which case the the system default will be used.

Reflow Text

By checking this option, reformatting text when replying to a message is on as default. Even if this option is on, reflow can still be turned off when replying to a message.

Quoting

This cycle gadget defines what type of quoting to apply to messages you reply to. Six different types are supported:

Use Global

This will use the quote type defined in the system type (or the system quoting if started from the Conference Configuration).

Default

This will put a ">" in front of all quoted lines.

None

No quoting will be applied to the message.

Def+Space

Will put a ">" in front of all lines and puts an extra space after the ">" on lines which has not been quoted before. Readability will be higher and it looks better.

Custom

Lets you enter your own character(s) to put before lines which are quoted in the string gadget to the right.

Initials

Messages will be quoted with the initials of the person that wrote the message replied to. If "Petter Nilsen" wrote a message your are replying to, "PN|" will be put in front of all quoted lines. This is commonly used on BBSes and should not be used on the Internet.

Reply String

This string will be added at the top of the message when you reply. It may contain several codes that is replaced with various information:

       %S        -   Subject
       %A        -   Sender address
       %N        -   Sender full name
       %F        -   Sender first name
       %L        -   Sender last name
       %R        -   Reciever address
       %E        -   Receiver name
       %C        -   Conference
       %I        -   MsgID
       %O        -   Original messagenr on remote system
       %D        -   Date

NOTE: Not all fields are available on all systems. If a field is not available, it will be left empty. A cycle gadget here may be used to tell THOR when to add a reply string:

          Use Global

The globally defined setting for reply string will be used on this system/conference (default).

          Only if moved

The reply string will only be used if the message in not replyed to in the same conference or to the same user.

          Always

The reply string will always be used when making a reply.

          Don't use

Don’t use any reply string on this system at all.

Type specific options

If this gadget is not ghosted, selecting it will start an external configuration program for the system type of the selected system.

3 Global Configuration

Use (config)

Lets you test the global configuration selected.

Save (config)

Saves the global configuration to disk.

Last Saved

Loads the last saved global configuration from disk. It will be used when you select Use (config).

Cancel

Cancels all changes made to the global configuration.

Global database options

This window is used for setting of parameters related to the message database.

Buffers

This value defines the number of buffers the buffering system uses. Note that 13 is the default number of buffers and you will not notice any significant speed difference by increasing this number, although reading and parsing messages in reference mode is slightly faster with a higher setting. The Buffer size has much greater influence on performance.

Buffer Size

This value is the size of each buffer. Default is 5096, and should not be any less (since this is the size of one of the hash-tables.) However, if you have the memory to spare, you should increase this for higher performance. These buffers are used by BBSRead.library and will be freed when bbsread.library is flushed out of memory or after a 5 minutes timeout if the library is not in use anymore.. Defining too large buffer here might make THOR slower because of overhead. Usually, 20000 is a good size of the buffer.

Msgs to Keep

This value is used when packing conferences. If the number is set to e.g. 100, all messages except for the newest 100 will be deleted when you pack messages This is the default for all conferences if nothing else is defined in the Chapter 4 [Conference Configuration], page 12. If you set the checkbox gadget to "Ignore", no messages will be packed with this parameter.

Days to Keep

This value is used when packing conferences, unless another value is defined in the Chapter 4 [Conference Configuration], page 12. When packing, all messages older than this number of days will be removed. This works seemless together with the Msgs to Keep value. If you set the checkbox gadget to "Ignore", no messages will be packed with this parameter.

Xpk Method

This is the Xpk method that will be used when packing messages, e.g. "NUKE" or "RAKE".

Auto Xpk Packing

If this gadget is checked, Xpk packing will be preformed when the messages are imported into the database. Otherwise, if will only happen when packing messages.

Haze Levels

This feature is named after Dave ’hazy’ Haynie. On emphasize/kills, any matching message can be given a special Haze Level . The various haze-levels are defined in the Global Configuration in ConfigTHOR and defines the minimum number of days the message is to remain in the messages database before it can be purged. Tip: If you want to keep your own messages (the ones you write yourself) in the database somewhat longer, setup a Mark Unread trigger on your own email address, then set a haze level that’ll keep the message in the database longer than other messages.

Global message options

Editor

This cycle gadget may be either set to "Internal", in which case the internal editor is used, or "External" which will activate the string gadget where you can type in the name and command line options to the external editor of your choice. Note that if you use an external editor that can open on a public screen, in is suggested that you use THOR’s public screen. It has the name "Thor.n", where n is 1 if you are running one copy of THOR. The following codes can be used on the command line for the external editor definition and will be replaced with the appropriate data by THOR:

%P - pubscreen name Note that this code will be at the end of the line passed to the editor.

Tab Size

This is the TAB size used in the internal editor and in the message window.

Auto Next Msg

This flag can be set if you want to go directly to the next message when you have replied to the current one.

No "Next Conf?" Requester

This flag can be set if you don’t want a window to pop up when you are finished reading the messages in a conference. THOR will instead automatically proceed to the next conference.

Auto Show First Message

If this option is selected, THOR will automatically show the first message of a conference when you enter it. If this is not selected, you will have to proceed to the "next" message to get THOR to show you the first in a conference.

Delayed Msg Unmark

When this option is selected, THOR will not mark the message you read as read until you proceed to the next message.

PGP Command

This is the full path and name of the PGP command you want to use when encrypting messages. If nothing is entered here, THOR will use the path defined in the envvariable PGPPATH (if available). Please refer to the PGP documentation for further information.

PGP SignID

This is the signature ID you want to use when PGP signing messages. "*" is the default and will use the first item in your secret key ring (which usually only contain one secret key: yours). Please refer to the PGP documentation for further information.

Reply String

This string will be added on top of messages when you reply. I may contain several codes that THOR will replace with various information: %S - Subject %A - Sender address %N - Sender name %F - Sender first name %L - Sender last name %R - Reciever address %E - Receiver name %C - Conference %I - MsgID %O - Original messagenr on remote system %D - Date NOTE: Not all fields are available on all systems. If a field is not available, it will be left empty. A cycle gadget is available that can be used to tell THOR when to add a reply string:

Don't use

Don’t use any reply string on this system at all.

Only if moved

The reply string will only be used if the message is not replied to in the same conference or to the same user.

Always

The reply string will always be used when making a reply.

Tag file

This is the tag file. A tag is a random joke or a cookie added at the end of messages you write. The file must have one tag on each line. If you want to use multiline tags, this can be done by using backslash n (\n) to separate lines in the tag. The field can be left empty, in which case no tag will be used. A cycle gadget lets you choose not to use a tag file at all or use the tag file you define here.

Signature

This is the global signature that is appended at the end of the message file before it is imported into the editor (and in front of any taglines). In a multi-line signature each line can be separated with a "\n" (backslash n). There are 3 methods available to add a signature:

String

The contents of the signature string gadget will contain the string to be imported into the message.

File

The contents of the signature string gadget is a file name of a text file that will be imported into the message.

Command

The contents of the signature string gadgets is a shell command that will receive arguments with the template "BBSNAME/K,CONFNAME/K" when run and must write the output to stdout.

A cycle gadget lets you choose to don’t use a signature on at all or use the signature you define here.

Global path options

This window is for different paths which THOR will use in its operation. The following paths can be defined here:

Save dir

With this option you can select a directory to be used as default in the file requester when you choose Save from the menu.

Download dir

This is the download directory where grabs, packets or mailboxes that you download from systems are searched for. Should be configured the same as in your communication program. If you collect mail using UUCP or SMTPd, this should normally be set to UUMAIL:

Upload dir

This is the upload directory where reply packets will be placed.

ARexx dir

This is the ARexx directory where THOR defaults to look for new ARexx scripts when adding macros. This directory is only a default for the file requester, you can use arexx scripts from other directories as well.

Temp dir

If configured, this directory will be used as temporary storage when packing/purging the message database.

DOS Output File

This is a CON: definitions that will be used for parsers and packers that outputs to stdout (command line). Normally, this should be left alone. The default setting is:

     "CON:0/11/640/200/THOR<->DOS Output/AUTO/SCREEN %s".

Global archivers

This window is for editing and adding of archivers. An archiver is used to uncompress grabs or message packets downloaded, and to compress replies to be uploaded. Several archivers are predefined and, in most cases, you don’t need to make any changes. However, these functions are available:

Arc Types

This is a list of available archiver types from which you may select one to edit or delete. Note that these names are only names, and do not actually have to be the command used.

Edit

Editing of the selected archiver’s parameters: "Archive Pattern" is the pattern used to identify files packed with this type. This pattern can be either the file name (e.g. "#?.lha") or the first bytes of the file, something like "$????2d6c68". This pattern is compared with the file in question, to check if they match. The"$" is used to define that this is a hex-pattern and the "?" is for each byte that is overlooked and not used. (In the example, the last "2d6c68" is the actual hex-pattern that defines this file to be a file archived with LhA.) "UnArchive Command" is the actual command used to unpack these archives. "Archive Command" is the actual command used to pack message packets that are to be uploaded to a system. Which packer to be used is defined in the Chapter 2 [Systems Configuration], page 1.

New

Adding of new archiver types to the list. You have to edit this new type afterwards by entering pattern, unarchive command, etc.

Delete

Delete the selected archiver type.

Global screen options

Msg Font

This option brings up a Font Requester. Only fixed-width fonts are shown here, and maximum size is set to 20 points. This font is used in the listview for message text. Using one of the fonts included with THOR is recommended, but any fixed-width font can be used.

List Font

Lets you choose which font to use in all other listviews except the message text.

FSE Font

Lets you choose the font used in the internal editor. Note: If this font is too big to fit into the window for the FSE, it will default to topaz/8.

Gadget Font

Lets you choose which font to used in all the button gadgets. If no font is selected, it will default to the Workbench screen font.

Screen Mode

Lets you choose Screen Mode, resolution and number of colors. Note: Selecting 8 colors means the GUI in THOR will look better. If you want ANSI or MIME attachments to look really good, 16 colors or higher is recommended. (Of course this is a tradeoff between speed/chipmem usage and ultracoolness). Please consider the number of colors used in any background pictures when selecting the number of colors for the screen here.

Palette

Lets you modify the palette, in 24 bits if supported by your hardware. Only the pens that THOR uses can be edited under OS 3.0.

Hide main window button panel

Use this option if you don’t want to have the bottom gadget row in the Main Window. The shortcuts for these gadgets will still work, but more space is used for message text.

Global ARexx scripts

THOR Startup

This is the ARexx script THOR will run async just before the Startup Window is opened and when THOR is uniconified.

THOR Quit

This is the ARexx script THOR will run sync just before any windows are closed when you quit THOR and when THOR is iconified.

Entering Systems

This is the ARexx script THOR will run sync when entering a system just before the Main window is opened, after parsing of any mail packets.

Leaving Systems

This is the ARexx script THOR will run sync when leaving a system, just before any windows are closed (and after any conference leaving script has been run).

Entering Confs

This is the ARexx script THOR will run sync when entering a conference, just after the first message in the conference has been shown. NOTE: This script will not be run when using the ARexx commando SHOWCONFERENCE to avoid any loop situation.

Leaving Confs

This is the ARexx script THOR will run sync when leaving a conference.

When the THOR Startup ARexx script is run async , THOR will proceed as usual and process ARexx commands at the same time.

When the other ARexx scripts are run sync , THOR will only prosess incoming ARexx commands and messages, and nothing else will be done. THOR will not proceed until all outstanding ARexx scripts have been completed.

4 Conference Configuration

This gadget lets you configure selected conferences in several ways.

Delete

Deletes the conference selected in the list of conferences. Note: Selecting this option will delete ALL messages and settings for this conference.

Sort

This option sorts the conferences alphabetically. Please note that the changes will not be saved to disk until you select "Save".

Rename

Renames the currently selected conference.

Top/Up/Down/Bottom

These buttons lets you rearrange the order of the conferences. Please note that the changes will not be saved to disk before you select "Save".

Days to Keep (conf)

All conferences will use the global option as default. However, if you want to enter a specific value for the chosen conferences, this will let you enter the number of days to keep messages. The option "Use This" must be selected in the cycle gadget for this value to have any effect.

Msgs to Keep (conf)

Lets you enter the maximum number of messages allowed in the selected conference. The option "Use This" must be selected from the cycle gadget for this value to have any effect.

Save (conf)

Saves changes made to all conferences. After rearranging conferences in the list, you must use this option to save the new conference order.

Cancel (conf)

Cancels all changes made to the conferences.

Please see [User Info], page 2, [Files/Paths], page 2 and [Import/Export], page 3 for the other options in this window.

Index 13

Index

7

7-bit Mail. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

A

About . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Add users. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Address. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Alias . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Ansi Menus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Arc Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Archivers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 ARexx dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Arexx script . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ARexx Scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Auto DL Private . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Auto Log Off . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Auto Next Msg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Auto Show First Message . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Auto Xpk Packing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

B

Bottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Buffer Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Buffers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

C

Cancel (conf) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Cancel (config). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Charset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Command Signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Conference Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Conference enter Arexx script . . . . . . . . . . . . . . . . . . . . . . . 2 Conference leave Arexx script . . . . . . . . . . . . . . . . . . . . . . . 2 Country . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

D

Days to Keep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Days to Keep (conf) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Delayed Msg Unmark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Delete Conference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Down. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Download dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Download directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

E

Edit Arc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 EMail address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Entering Confs Arexx Script . . . . . . . . . . . . . . . . . . . . . . . 11 Entering Systems Arexx Script . . . . . . . . . . . . . . . . . . . . . 11 Event Arc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

F

File Signature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 FSE Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

G

Gadget Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Get Bulletins . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Get Newfiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Global Configuration. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Global Paths and Files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Grabname . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

H

Haze Levels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Hide main window button panel. . . . . . . . . . . . . . . . . . . . 11

I

In directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

L

Last Saved (config) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Leaving Confs Arexx Script . . . . . . . . . . . . . . . . . . . . . . . . 11 Leaving Systems Arexx Script. . . . . . . . . . . . . . . . . . . . . . 11 Line Length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 List Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

M

Mailpacket name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Mark Own Msgs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 MIME . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Msg Font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Msgs to Keep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 12

N

No "Next Conf?" Requester. . . . . . . . . . . . . . . . . . . . . . . . . 7

O

Out directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

Palette . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 PGP Command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 PGP SignID . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Phone number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Pictures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

Q

Quote Char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Quoted-Printable encoding . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Index 14

Quoting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

R

Reflow Text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Rename conference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Reply String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 7

S

Save (conf) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Save (config) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Save dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Screen Mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Screen Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Screen Picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Script Flags. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Signature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2, 7 Signature File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Sort conferences . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Street Address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 String Signature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 System enter Arexx script. . . . . . . . . . . . . . . . . . . . . . . . . . . 2 System leave Arexx script . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 System type options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Systems Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

T

Tab Size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Tag file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 9 Tagfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Temp dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 THOR Quit Arexx Script . . . . . . . . . . . . . . . . . . . . . . . . . . 11 THOR Startup Arexx Script . . . . . . . . . . . . . . . . . . . . . . . 11 Top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Type options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

U

Up . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 Upload dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Upload directory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Use (config). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Use Colors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 User name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

W

Window Picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

X

Xpk Method . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3, 6

Table of Contents

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 About . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Systems Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . 1 User information. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Files and paths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 Import and Export options . . . . . . . . . . . . . . . . . . . . . . . . . 3 Type specific options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3 Global Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Global database options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Global message options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Global path options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Global archivers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 Global screen options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Global ARexx scripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4 Conference Configuration . . . . . . . . . . . . . . . . . . . . . . 12 Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13